If lim x tends to 0 , (cos x +a sin bx)^1/x = e^2 , prove that ab=2

Grade:11

1 Answers

NANDA
9 years ago
lim X->0 (1+x)^1/x =e^1 (1+asinbx)^1/x=e^k k=limx->0 asinbx/x=ab k=2(given) hence,ab=2